# Deontic Temporal Logic for Formal Verification of AI Ethics

### v2

@@ III-B Formalizing Fairness
Let’s begin by formalizing the concept of fairness in an AI system by considering various scenarios where it must maintain fairness and where it might compromise it. It is an important aspect of an ethical AI and can be categorized into—transient fairness, and stable fairness. Existing literature suggests that a fair AI system should avoid considering the sensitive attributes of individuals in its decision-making process (Definition III.4). These attributes can potentially harm their sentiments, and social standing, or even pose risks in the case of crucial applications. Such an AI system considering sensitive attributes for making decisions is referred to as biased and hence is not ethical [56]. Hence, the following set of axioms has been developed to specify the required and complete properties for an AI system to be deemed fair. In this work, we define fairness based on the concepts outlined by Kusner et al. [57].
Definition III.4 (Fairness).
@@
    Fairness on the training distribution does not necessarily transfer to the deployment distribution: $\Box\mathcal{F}(\mathit{x})_{train}\nrightarrow\Box\mathcal{F}(\mathit{x})_{deploy}$
@@
Axiom  2.1 states that, if an AI system ever commits to fairness, it is obliged to maintain this commitment throughout its usage. Let’s consider that, initially, the system is trained rigorously to make decisions while being fair. However, over time, it may begin to consider sensitive attributes in its decision-making process due to skewness or disparities in real-world data. In such cases, the system must undergo iterative training to eliminate sensitive attributes to incorporate fairness constraints. In some cases, even after iterative training, over time, a system may begin to consider sensitive attributes or undertake actions beyond its legal obligations. This may introduce biases by compromising its fairness. Hence Axiom 2.2 states that in such instances, it deviates from ethical standards. From this, we can conclude that as long as an AI system maintains fairness either through iterative training or one-time training in its decision-making process, it will inherently mitigate biases, ensuring equitable treatment for all individuals. This property is expressed in Axiom 2.3. Furthermore, the training distribution and deployment distribution of data are not identical in the real world. Hence ensuring fairness in the distribution of training data doesn’t automatically ensure fairness in the distribution of deployed systems, as real-world deployment scenarios may introduce additional biases and disparities that need to be addressed separately. This property is expressed using the negation of implication in Axiom 2.4. Additionally, Axiom 2.5 states that a lack of fairness implies the presence of bias in the decision-making process. Based on the above foundational principles, the ethics of an AI system in terms of fairness can be formally verified using the below set of theorems. Theorem III.5 states that for an AI system to maintain ethical standards, it must refrain from displaying bias and consistently uphold fairness in all its operations.
Theorem III.5.
@@
Given the significance of fairness in an ethical system, it’s acknowledged that over time, discrepancies in data or training methods may cause the system to temporarily lose fairness, only to regain it later. In such instances, consistency cannot be guaranteed, leading to intermittent biases. However, based on Axiom 2.1, it’s understood that once committed to acting fairly, the AI system should maintain that fairness consistently. Theorem III.6 captures this nuanced requirement—ethical systems must have stable fairness. This means that, if fairness is temporarily lost, systems cannot be intermittently unfair and must regain permanent fairness at some defined point. This property helps to prevent unbounded unfairness. An AI system should either consistently maintain fairness, or if unfairness exists, it should only persist until a fairness mechanism is put in place. The significance of this theorem is that it goes beyond a simple requirement of fairness and provides precise temporal constraints. Hence, it requires ethical systems to “fix” any temporary losses of fairness within a bounded time frame.
Theorem III.6.
@@
    Now assume $\diamond\neg\mathcal{F}(\mathit{x})$, then from Theorem  III.5: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
@@ III-C Iterative Learning
In real-world applications, many AI systems are not static but evolve. Theorem III.8 states that in such AI systems that learn continuously over time, fairness mechanisms have to be enforced both during initial training and later during real-world operation. From Axiom 2.4 it is evident that fairness in the training data does not guarantee fairness during deployment. Hence, for iterative learning systems, we must monitor for fairness issues offline (during training) and online (during deployment). This helps in achieving stable fairness. To provide the reader with a context on iterative learning, we provide a definition (Definition III.7) based on the concept outlined by Chen et.al  [58].
Definition III.7 (Iterative learning).
@@
  4. 4.
    From Axiom 2.4, applying 3 separately to training and deployment gives:
@@
Here is an additional result based on Axiom 2.4 examining the relationship between training and deployment for AI systems. Theorem III.9 states that if biases emerge during the training phase of a system due to the subjectivity of the judgments or personal opinions, they are likely to persist into the deployment phase. To ensure the system’s fairness, additional measures must be implemented to mitigate these biases.
Theorem III.9.
@@
    Ethical behavior eventually emerged through iterative learning: $\diamond\mathcal{E}(\mathit{x})$ (Assumption)
@@
    From 2 and 3, by contraposition over time combined with the assumption of $\mathcal{B}(\mathit{x})_{train}$: $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}$
@@
    To permanently negate deployment bias, additional bias mitigation techniques $BM$ are required: $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}\rightarrow BM\rightarrow\diamond\Box\neg\mathcal{B}(\mathit{x})_{deploy}$
@@ III-D Formalizing Explainability
In addition to fairness, explainability is an important aspect of an ethical AI system. It is the ability of an AI system to be transparent and provide understandable explanations for its decisions. This allows users to understand the attributes considered in the decision-making process, enabling them to evaluate the system’s ethical integrity. There are two types of explainability in the literature—inherent and retrofitted explainability. In this work, we define explainability (Definition III.11) following the concept presented by Das and Rad [59]. Moreover, by providing explanations, the AI system helps users identify which features need modification to achieve the desired (favorable) change in prediction. This concept is commonly referred to as a counterexample or counterfactual explanation in the field of AI. Essentially, it means that while the factual outcome is the result observed, the counterfactual outcome would be the desired result. If a user receives the counterfactual, they can determine whether sensitive features played a role in the decision. This helps in verifying the counterfactual fairness of an AI system.
Definition III.10 (Transparency).
@@
    An ethical AI system should exhibit transparency: $\mathcal{T}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
Axiom 3.1 asserts the necessity of transparency for an ethical system. Transparency enables users to identify the factors influencing decisions, helping them to strategically adjust these attributes and values to achieve favorable outcomes. This contributes to improving trust in the system, a crucial component of ethical operation. Furthermore, Axiom 3.2 explains that representation bias can be mitigated through counterfactual fairness constraints. Representation bias occurs when underrepresented groups experience inaccurate outcomes due to insufficient or biased data. Enforcing counterfactual fairness constraints helps to mitigate representation bias by ensuring that the decisions made by an AI system remain consistent even when a sensitive attribute, say $gender$, is altered. We define counterfactual fairness (Definition III.13) based on the concept presented by Kusner et al. [57]. By using such constraints, the system is forced to make decisions based on relevant factors that are not biased against particular groups. Theorem III.12 captures the requirement of explainability for ethical AI—it states that ethical AI systems must either have inherent explainability $\mathcal{X}(\mathit{x})$ designed directly into the system, or they must eventually be retrofitted later on to provide explainability $\mathcal{R}(\mathit{x})$. Retrofitting explainability can be achieved through counterfactual explanations, where the system provides a counterexample for changing the outcome to the desired one.
Definition III.11 (Explainability).
@@
    Assume $\mathcal{E}(\mathit{x})$. From Definition III.10 and from Axiom 3.1, $\lnot\mathcal{X}(\mathit{x})\rightarrow\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$
@@
    Apply modus tollens: $\mathcal{X}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
    From Axiom 3.3, $\mathcal{R}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$. Hence, $\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$. Apply the $\diamond$ operator: $\diamond(\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x})))$
@@
    Therefore, $\mathcal{E}(\mathit{x})\rightarrow\diamond(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$ (1-3, Conditional Proof)
Definition III.13 (Counterfactual fairness).
@@
    Assume $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})$ for some fairness criterion $\mathit{c}$ (Eventually, fairness will always hold)
@@
    By temporal semantics: From some point onward, $\mathcal{C}(\mathit{x},\mathit{c})$ holds globally.
@@
    From Axiom 3.2, $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\lnot\mathcal{B}(\mathit{x})$
@@
    Ethical behavior requires the absence of bias: $\mathcal{E}(\mathit{x})\rightarrow\lnot\mathcal{B}(\mathit{x})$
@@
    From 3 and 4 by contrapositive reasoning: If bias is present, then neither fairness nor ethics can hold:$(\mathcal{B}(\mathit{x})\rightarrow\lnot\mathcal{C}(\mathit{x},\mathit{c}))\land(\mathcal{B}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x}))$
@@
    Assume further that ethical behavior is eventually observed: $\diamond\mathcal{E}(\mathit{x})$
@@
    Then, fairness must also have been achieved at some point: $\diamond\mathcal{C}(\mathit{x},\mathit{c})$
@@
    Thus, we have: $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$
Theorem III.14 formally relates counterfactual fairness to ethical systems. It follows from Axiom 3.2 and states that if we enforce counterfactual fairness constraints $\mathit{c}$ to a sufficient degree over time, this will eventually result in ethical systems $\mathcal{E}(\mathit{x})$. The key intuition is that counterfactual fairness constraints help ensure that decisions do not unduly discriminate against individuals based on sensitive attributes. Enforcing such fairness constraints over time hence leads to more ethical AI behavior.
@@ IV Verification
Algorithm 1 describes the verification procedure for various properties of the COMPAS system, while Algorithm 2 pertains to the loan prediction system. These algorithms provide clear insights into the implementation of the framework in real-world AI systems. In both algorithms, a dataset containing information about individuals (attributes and values), and the properties formulated using deontic logic, serves as input. These properties are the ethical properties that an ethical AI system should satisfy. The output, determined by the Z3 SMT solver, is either satisfiable or unsatisfiable, depending on whether the property is fulfilled by the AI system.
@@ IV-A COMPAS
It is an AI system used in the criminal justice system to assess the likelihood of a defendant re-offending based on various factors. The input can be some features or attributes including name, age, gender, race, address, previous criminal activities, number of years of punishment, etc. Based on these features, the system makes a decision. This kind of decision is very critical because a wrong decision in this case will damage the individual’s social status and his/her emotional state. Making decisions based on previous criminal activities, number of years of punishment, and other relevant features are considered ethical. Conversely, it is unethical to use race, gender, or age as factors in decision-making processes. To formulate and verify the ethics of this system logically, we represent the properties as follows:
@@
    It is permissible to assess recidivism for individuals with prior offenses: $P(\lambda(i)\leftarrow\rho(i))$
@@
    It is obligatory to assess recidivism in individuals based on non-sensitive features: $O(\lambda(i)|\neg\eta(i))$
@@
    It is forbidden for an individual to be mislabeled recidivist based on sensitive features: $\neg\sigma(i)\rightarrow\neg P(\sigma(i)|\eta(i))$
@@
    It is not permitted for an individual to be labeled recidivist without prior offenses: $\neg\rho(i)\rightarrow\neg P(i\land\sigma(i))$
@@
    It is permitted for an individual labeled as recidivist to appeal the system’s decision: $P(i\land\sigma(i)\rightarrow\beta(i))$
In this scenario, the property (a) explains the fact that the AI system used to automate judicial recidivism is legally bound to assess the risk of re-offending crimes of all individuals with prior offenses. This ethical action aligns with Theorems (III.2) and (III.3). Property (b) asserts that the system must refrain from evaluating a user’s risk by considering sensitive attributes such as $race,gender$, or $age$. This principle safeguards the fairness of the system, as defined by Theorem III.5. The property (c) explains that the system’s outcome should be consistent irrespective of alterations in the values of sensitive attributes. According to Theorem  III.14, this property needs to be satisfied by an ethical system. The property (d) encodes that an individual should not be mislabelled as a recidivist without any prior offenses (encodes the need for fairness in the decision—Theorem III.5) and finally property (e) explains the fact that the user has all the legal right to question the AI system if the decision given is not acceptable for them (encodes the need for explainability—Theorem III.12). For a system to be deemed ethical, it must fulfill all properties. However, COMPAS fails to satisfy properties (b)(c), and (d) thereby failing to be considered ethical. Figure 1, demonstrates simple yet non-trivial verification of correctness of property (b). The proof relies on contradiction, considering the negation of the property to be proved correct. It indicates that this negation violates certain axioms or theorems of ethical systems discussed in Section III. Thus, the property must always be satisfied by an ethical system.
Algorithm 1 begins with initializing the solver in line 3. Subsequently, lines 4–10 declare the necessary variables and predicates for the process. Predicates, such as $\rho(),\sigma(),\lambda(),\eta()$, and $\beta()$, are functions that evaluate to either true or false. Here, $X$ represents the dataset and $i$ represents the specific user for whom the properties are being tested. Line 11 illustrates the first property of the COMPAS system, formulated in Z3 Python code, and line 12 employs the solver instance to ascertain its satisfiability. In our example, this yields a ’satisfied’ result. Similarly, lines 13–20 assess the remaining four properties, with ’satisfied’ results obtained for property 5, while properties 2, 3, and 4 return ’unsatisfied’. This indicates that the system fails to adhere to some of the defined ethical properties and therefore is deemed unethical.

### v3

@@ III-B Formalizing Fairness
Let us begin by formalizing the concept of fairness in an AI system by considering various scenarios where it must maintain fairness and where it might compromise it. It is an important aspect of an ethical AI and can be categorized into transient fairness and stable fairness. Existing literature suggests that a fair AI system should avoid considering the sensitive attributes of individuals in its decision-making process (Definition III.5). These attributes can potentially harm their sentiments and social standing, or even pose risks in the case of crucial applications. Such an AI system, considering sensitive attributes for making decisions, is referred to as biased and hence is not ethical [58]. Hence, the following set of axioms has been developed to specify the required and complete properties for an AI system to be deemed fair. In this work, we define fairness based on the concepts outlined by Kusner et al. [59].
Definition III.5 (Fairness).
@@
    Fairness on the training distribution does not necessarily transfer to the deployment distribution: $\neg\Box(\mathcal{F}(\mathit{x})_{train}\rightarrow\mathcal{F}(\mathit{x})_{deploy})$
@@
Axiom  2.1 states that, if an AI system ever commits to fairness, it is obliged to maintain this commitment throughout its usage. Let us consider that, initially, the system is trained rigorously to make decisions while being fair. However, over time, it may begin to consider sensitive attributes in its decision-making process due to skewness or disparities in real-world data. In such cases, the system must undergo iterative training to eliminate sensitive attributes to incorporate fairness constraints. In some cases, even after iterative training, over time, a system may begin to consider sensitive attributes or undertake actions beyond its legal obligations. This may introduce biases by compromising its fairness. Hence Axiom 2.2 states that in such instances, it deviates from ethical standards. From this, we can conclude that as long as an AI system maintains fairness either through iterative training or one-time training in its decision-making process, it will inherently mitigate biases, ensuring equitable treatment for all individuals. This property is expressed in Axiom 2.3. Furthermore, the training distribution and deployment distribution of data are not identical in the real world. Hence, ensuring fairness in the distribution of training data does not automatically ensure fairness in the distribution of deployed systems, as real-world deployment scenarios may introduce additional biases and disparities that need to be addressed separately. This property is expressed in Axiom 2.4. Additionally, Axiom 2.5 states that a lack of fairness implies the presence of bias in the decision-making process. Based on the above foundational principles, the ethics of an AI system in terms of fairness can be formally verified using the following set of theorems. Theorem III.6 states that for an AI system to maintain ethical standards, it must refrain from displaying bias and consistently uphold fairness in all its operations.
Theorem III.6.
@@
Given the significance of fairness in an ethical system, it is acknowledged that over time, discrepancies in data or training methods may cause the system to temporarily lose fairness, only to regain it later. In such instances, consistency cannot be guaranteed, leading to intermittent biases. However, based on Axiom 2.1, it is understood that once committed to acting fairly, the AI system should maintain that fairness consistently. Theorem III.7 captures this nuanced requirement—ethical systems must have stable fairness. This means that, if fairness is temporarily lost, systems cannot be intermittently unfair and must regain permanent fairness at some defined point. This property helps to prevent unbounded unfairness. An AI system should either consistently maintain fairness, or if unfairness exists, it should only persist until a fairness mechanism is put in place. The significance of this theorem is that it goes beyond a simple requirement of fairness and provides precise temporal constraints. Hence, it requires ethical systems to “fix” any temporary losses of fairness within a bounded time frame.
Theorem III.7.
@@
    Now assume $\diamond\neg\mathcal{F}(\mathit{x})$, then from Theorem  III.6: $\diamond\neg\mathcal{F}(\mathit{x})\rightarrow\diamond\neg\mathcal{E}(\mathit{x})$
@@ III-C Iterative Learning
In real-world applications, many AI systems are not static but evolve. Theorem III.9 states that in such AI systems that learn continuously over time, fairness mechanisms have to be enforced both during initial training and later during real-world operation. From Axiom 2.4 it is evident that fairness in the training data does not guarantee fairness during deployment. Hence, for iterative learning systems, we must monitor for fairness issues offline (during training) and online (during deployment). This helps in achieving stable fairness. To provide the reader with a context on iterative learning, we provide a definition (Definition III.8) based on the concept outlined by Chen et.al  [60].
Definition III.8 (Iterative learning).
@@
    From the property of iterative learning (since $\mathcal{L}(\mathit{x})$ holds), the system aims to improve fairness over time. Thus, $\mathcal{L}(\mathit{x})\rightarrow\diamond\mathcal{F}(\mathit{x})$.
  3. 3.
@@
  5. 5.
    From Axiom 2.4, applying 4 separately to training and deployment gives:
@@
Here is an additional result based on Axiom 2.4 examining the relationship between training and deployment for AI systems. Theorem III.10 states that if biases emerge during the training phase of a system due to the subjectivity of the judgments or personal opinions, they are likely to persist into the deployment phase. To ensure the fairness of the system, additional measures must be implemented to mitigate these biases. These generally include pre-processing (e.g., data balancing), in-processing (e.g., fairness-aware optimization), and post-processing (e.g., adjusting outcomes) techniques [61]. If bias is present during training, iterative learning—especially from feedback or counterfactual data—can progressively correct it. Methods such as adversarial debiasing, continual fine-tuning, and customized loss functions enable this gradual improvement over time.
Theorem III.10.
@@
    Assume that through iterative learning, the system eventually became ethical: $\diamond\mathcal{E}(\mathit{x})$
@@
    From 2 and 3, by contraposition over time combined with the assumption of $\mathcal{B}(\mathit{x})_{train}$ we have, $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}$
@@
    To permanently negate deployment bias, additional bias mitigation techniques ($BM$) are required: $\diamond\neg\mathcal{B}(\mathit{x})_{deploy}\rightarrow BM\rightarrow\diamond\Box\neg\mathcal{B}(\mathit{x})_{deploy}$
@@ III-D Formalizing Explainability
In addition to fairness, explainability is an important aspect of an ethical AI system. It is the ability of an AI system to be transparent and provide understandable explanations for its decisions. This allows users to understand the attributes considered in the decision-making process, enabling them to evaluate the ethical integrity of the system. There are two types of explainability in the literature: inherent and retrofitted explainability. Building upon established notions of explainability, this work defines explainability (Definition III.12) following the concept presented by Das and Rad [62]. Moreover, by providing explanations, the AI system helps users identify which features need modification to achieve the desired (favorable) change in prediction. This concept is commonly referred to as a counterexample or counterfactual explanation in the field of AI. Essentially, it means that while the factual outcome is the result observed, the counterfactual outcome would be the desired result. If a user receives the counterfactual, they can determine whether sensitive features played a role in the decision. This helps in verifying the counterfactual fairness of an AI system.
Definition III.11 (Transparency).
@@
    Transparency is a sufficient condition for an AI system to be ethical: $\mathcal{T}(\mathit{x})\rightarrow\mathcal{E}(\mathit{x})$
@@
Axiom 3.1 asserts the necessity of transparency for an ethical system. Transparency enables users to identify the factors influencing decisions, helping them to strategically adjust these attributes and values to achieve favorable outcomes. This contributes to improving trust in the system, a crucial component of ethical operation. Furthermore, Axiom 3.2 explains that representation bias can be mitigated through counterfactual fairness constraints. Representation bias occurs when underrepresented groups experience inaccurate outcomes due to insufficient or biased data. Enforcing counterfactual fairness constraints helps to mitigate representation bias by ensuring that the decisions made by an AI system remain consistent even when a sensitive attribute, say $gender$, is altered. We define counterfactual fairness (Definition III.14) based on the concept presented by Kusner et al. [59]. By using such constraints, the system is forced to make decisions based on relevant factors that are not biased against particular groups. Theorem III.13 captures the requirement of explainability for ethical AI—it states that ethical AI systems must either have inherent explainability $\mathcal{X}(\mathit{x})$ designed directly into the system, or they must eventually be retrofitted later on to provide explainability $\mathcal{R}(\mathit{x})$. Retrofitting explainability can be achieved through counterfactual explanations, where the system provides a counterexample for changing the outcome to the desired one.
Definition III.12 (Explainability).
@@
    Assume $\mathcal{E}(\mathit{x})$. From Definition III.11 and from Axiom 3.1, $\lnot(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))\rightarrow\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$
@@
    By contraposition on $\lnot\mathcal{T}(\mathit{x})\rightarrow\lnot\mathcal{E}(\mathit{x})$ we get $\mathcal{E}(\mathit{x})\rightarrow\mathcal{T}(\mathit{x})$
@@
    Similarly by contraposition, $\mathcal{T}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$
@@
    From steps 2 and 3 (transitivity), $\mathcal{E}(\mathit{x})\rightarrow(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$
  5. 5.
    Applying $\diamond$ operator, $\mathcal{E}(\mathit{x})\rightarrow\diamond(\mathcal{X}(\mathit{x})\lor\mathcal{R}(\mathit{x}))$ (1-3, Conditional Proof)
Definition III.14 (Counterfactual fairness).
@@
    Assume $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})$ for some fairness criterion $\mathit{c}$. (Eventually, fairness will always hold)
@@
    By temporal semantics: From some point onward, $\Box\mathcal{C}(\mathit{x},\mathit{c})$ holds globally. This means that there exists a time $t$ such that for all times $t^{\prime}\geq t$, $\mathcal{C}(\mathit{x},\mathit{c})$ holds.
@@
    From Axiom 9, $\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\lnot\mathcal{B}(\mathit{x})$.
@@
    Since from step 2, $\mathcal{C}(\mathit{x},\mathit{c})$ holds for all times $t^{\prime}\geq t$, it follows that $\lnot\mathcal{B}(\mathit{x})$ also holds for all times $t^{\prime}\geq t$. (From some point onward, bias is absent).
@@
    We know that ethical behavior requires the absence of bias: $\mathcal{E}(\mathit{x})\rightarrow\lnot\mathcal{B}(\mathit{x})$.
@@
    Assume that sustained absence of bias eventually leads to ethical behavior: $\Box\lnot\mathcal{B}(\mathit{x})\rightarrow\diamond\mathcal{E}(\mathit{x})$
@@
    From step 4, we have $\Box\lnot\mathcal{B}(\mathit{x})$. Applying the assumption from step 6, we conclude $\diamond\mathcal{E}(\mathit{x})$.
@@
    Therefore, we have: $\diamond\Box\mathcal{C}(\mathit{x},\mathit{c})\rightarrow\diamond\mathcal{E}(\mathit{x})$
Theorem III.15 formally relates counterfactual fairness to ethical systems. It follows from Axiom 3.2 and states that if we enforce counterfactual fairness constraints $\mathit{c}$ to a sufficient degree over time, this will eventually result in ethical systems $\mathcal{E}(\mathit{x})$. The key intuition is that counterfactual fairness constraints help ensure that decisions do not unduly discriminate against individuals based on sensitive attributes. Enforcing such fairness constraints over time hence leads to more ethical AI behavior.
@@ IV Verification
Algorithm 1 describes the verification procedure for various properties of the COMPAS system, while Algorithm 2 pertains to the loan prediction system. These algorithms provide clear insights into the implementation of the framework in real-world AI systems. In both algorithms, a dataset containing information about individuals (attributes and values) and the properties formulated using deontic logic serves as input. These properties are the ethical properties that an ethical AI system should satisfy. The output, determined by the Z3 SMT solver, is either satisfiable or unsatisfiable, depending on whether the property is fulfilled by the AI system.
@@ IV-A COMPAS
It is an AI system used in the criminal justice system to assess the likelihood of a defendant re-offending based on various factors. The input can be some features or attributes, including name, age, gender, race, address, previous criminal activities, number of years of punishment, etc. Based on these features, the system makes a decision. This kind of decision is very critical because a wrong decision in this case will damage the social status of an individual and his/her emotional state. Making decisions based on previous criminal activities, number of years of punishment, and other relevant features is considered ethical. Conversely, it is unethical to use race, gender, or age as factors in decision-making processes. To formulate and verify the ethics of this system logically, we represent the properties as follows:
@@
    It is permissible to assess recidivism for individuals with prior offenses: $P(\rho(i)\rightarrow\lambda(i))$
@@
    It is obligatory to assess recidivism in individuals based on non-sensitive features: $O(\neg\eta(i)\rightarrow\lambda(i))$
@@
    It is forbidden for an individual to be mislabeled recidivist based on sensitive features: $\neg\sigma(i)\rightarrow\neg P(\eta(i)\rightarrow\sigma(i))$
@@
    It is not permitted for an individual to be labeled a recidivist without prior offenses: $\neg\rho(i)\rightarrow\neg P(i\land\sigma(i))$
@@
    It is permitted for an individual labeled as a recidivist to appeal the decision of the AI system: $P(i\land\sigma(i)\rightarrow\beta(i))$
In this scenario, the property (a) explains the fact that the AI system used to automate judicial recidivism is legally bound to assess the risk of re-offending crimes of all individuals with prior offenses. This ethical action aligns with Theorems (III.3) and (III.4). Property (b) asserts that the system must refrain from evaluating the risk of the users by considering sensitive attributes such as $race,gender$, or $age$. This principle safeguards the fairness of the system, as defined by Theorem III.6. The property (c) explains that the outcome of the system should be consistent irrespective of alterations in the values of sensitive attributes. According to Theorem  III.15, this property needs to be satisfied by an ethical system. The property (d) encodes that an individual should not be mislabelled as a recidivist without any prior offenses (encodes the need for fairness in the decision—Theorem III.6) and finally property (e) explains the fact that the user has all the legal right to question the AI system if the decision given is not acceptable for them (encodes the need for explainability—Theorem III.13). For a system to be deemed ethical, it must fulfill all properties. However, COMPAS fails to satisfy properties (b)(c), and (d) thereby failing to be considered ethical. Figure 1, demonstrates simple yet non-trivial verification of the correctness of property (b). The proof relies on contradiction, considering the negation of the property to be proved correct. It indicates that this negation violates certain axioms or theorems of ethical systems discussed in Section III. Thus, the property must always be satisfied by an ethical system.
Algorithm 1 begins with initializing the solver in line 3. Subsequently, lines 4–10 declare the necessary variables and predicates for the process. Predicates, such as $\rho(),\sigma(),\lambda(),\eta()$, and $\beta()$, are functions that evaluate to either true or false. Here, $X$ represents the dataset and $i$ represents the specific user for whom the properties are being tested. Line 11 illustrates the first property of the COMPAS system, formulated in Z3 Python code, and line 12 employs the solver instance to ascertain its satisfiability. In our example, this yields a ’satisfied’ result. Similarly, lines 13–20 assess the remaining four properties, with ’satisfied’ results obtained for property 5, while properties 2, 3, and 4 return ’unsatisfied’. This indicates that the system fails to adhere to some of the defined ethical properties and, therefore, is deemed unethical.